User Tools

Site Tools


operations:log

Operations

Schedules

Operations Meetings

Data Handling and Disk Logistics

Transfers

Please use the Transfers Page for information on data transfer.

Modules

See the Modules Page

AuScope IVS Module Allocation

If there are any questions or issues with module logistics, please contact auscope-operations@list.utas.edu.au with the details.

See Experiment and Module Catalog for module allocations, inventory, transfers, upcoming and historic experiments;

Telescope calibration/monitoring procedures and notes

Stairwell Displays

AuScope Operations Log

Please enter any significant events related to the AuScope VLBI project, including observations, configuration changes, hardware upgrades etc.

Katherine repairs

Three major repairs were finished on 2/12/15

* DBBC * IF wiring * Anemometer

DBBC: The Katherine DBBC is fully repaired as of 2/12/15. It originally failed 24/11/15, during r1715 with a total comms failure. This was eventually traced to the PCI7200 board, where the driver chip had burnt out. Replacing this board allowed Katherine to resume observations but with some apparent communication problems with the core2 units - bbc attenuations were being mis-set dur to a stuck bit in the communications line. The problem developed further after an initial inspection and no communications with either the CoMos or Core power levels could be made. The problem was eventually resolved on 1/12/15 when the PCI9111 board was replaced, but it may be that it was a connection issue instead - the PC frame is warped at Ke and it's possible to have some PCI lnes now connected through the backplane to the CPU.

IF wiring: The power level of SRCP going into CoMo D was unusually low in testing after the repair. It appears that the hardline or SMA connector was not making a good connection between the IF box and DBBC. Removing a replacing the hardline between the IF box output and DBBC radically improved the power level (~30 dB).

Anemometer: The anemometer at Katherine has suffered a fault and is reporting 1000 km/h winds. It has been turned off (using ispke) to avoid triggering spurious windstows. Until it is repaired, observers should use the weather radar at Katherine to look out for storm activity at Katherine.

2015/12/02 09:04 · Jamie McCallum · 0 Comments

Hb - S/X replaced and new pointing model

The S/X was replaced on Hb12 on 13/11/15. Initial tests with Y-factor tests indicate good X-band Tsys (~78 K) but fairly poor S-band Tsys (95-105 K). Pointing data was collected over the weekend and a new pointing model was generated (2015.317) but the corrections are small.

2015/11/16 04:51 · Jamie McCallum · 0 Comments

New IFBox servers for Hb, Ke and Yg

A new version of the IFBox server has been written and installed at Hb, Ke and Yg. It should be more robust than the previous version and reduce the number of spurious “odd response from the pic” messages. The new software is /usr2/AuscopeUtils/monitoring/PIC_interfaces/Monica.IFBox.Sep2015.py - the /etc/init.d/Monica.IFBox scripts have been updated to use this from now on.

2015/09/03 05:45 · 0 Comments

NTP servers updated on FS and Mk5 machines

The list of NTP servers has been updated on the field system and Mk5 computers. All have had the following servers added:

0.pool.ntp.org
1.pool.ntp.org
2.pool.ntp.org
3.pool.ntp.org

Procedure:

su
vi /etc/ntp.conf
(added the listed servers)
/etc/init.d/ntp restart

Computers updated:

  • mkv
  • pcfshb
  • pcfske
  • pcfsyg
  • mk5hb
  • mk5ke
  • mk5yg
  • pcfsho
  • mk5streamhb
  • mk5ho
  • mk5-2ke
  • mk5-2yg
  • hobart
2015/08/20 06:10 · 0 Comments

Log monitor updated

The log monitor has had a major update. New features:

  • The station name is more prominent
  • The name of the current scan is given
  • Current disk usage and data recorded in previous scan are given
  • A disk volume offset can be set. This is useful when comparing to predicted data volumes from the summary file in the case, for example of a module being changed part way through an experiment, or appending to a module with another experiment on it.
  • There's a test button for the SMS so you can check that text messages are getting through
  • For the Hobart 26m, the cryogenics pressure and 20K and 70K stage temperatures are reported. If the pressure exceeds 5 microns of Hg two times in a row, the alarm is sounded.
  • 26m cryo status monitoring can be turned off using a button under the “Configure” menu. This is useful if you know (and have reported) the problem but want to silence the alarms while it's being fixed.
  • GUI has been rearranged and fonds changed etc.

If you experience any problems with the new version, the old one can be used by typing pcfs_log_monitor_old.pl at the ops2 prompt.

2015/08/20 04:29 · 0 Comments
/home/www/auscope/opswiki/data/pages/operations/log.txt · Last modified: 2019/01/25 04:45 by Warren Hankey